OpenStudy (emilyjones284):

Please help ill give a medal...(: If f(x)=1/2x-6 , what is the equation for f^–1(x)?

OpenStudy (anonymous):

\[y = \frac{ 1 }{ 2x-6 }\] \[2x-6 = \frac{ 1 }{ y }\] \[x = \frac{ 1 }{ 2y } + 3\] Now interchange x with y for getting the inverse function.

OpenStudy (emilyjones284):

hmmm im still confused

OpenStudy (emilyjones284):

the beginning equation was \[f(x)=\frac{ 1 }{ 2 }x-6\]

OpenStudy (anonymous):

oh, i read it that way. Now all you need to do is bring 'x' on one side and leave all other terms the other side.

OpenStudy (emilyjones284):

what do you mean by interchanging x and y to get the inverse?

OpenStudy (anonymous):

In place of x,put y. In place of y,put x. From the definition of inverse, you can do this.

OpenStudy (emilyjones284):

ok thank you